Using the income statement information for Omega Corporation that follows, prepare a vertical analysis of the income statement for Omega Corporation. If required, round percentages to one decimal place. Percentage sign will appear automatically.

Omega Corporation

Income Statement

1

Amount

Percentage

2

Sales

$690,000.00

3

Cost of goods sold

372,600.00

4

Gross profit

$317,400.00
Answer #1

sing the income statement information for Omega Corporation that follows, prepare a vertical analysis of the income statement for Omega Corporation. If required, round percentages to one decimal place. Percentage sign will appear automatically.

Omega Corporation

Income Statement

1

Amount

Percentage

2

Sales

$690,000.00

690000/690000 = 100%

3

Cost of goods sold

372,600.00

372600/690000 = 54%

4

Gross profit

$317,400.00

317400/690000 = 46%

Cost of goods sold percentage = Cost of goods sold/Sales revenue

Gross profit percentage = Gross profit/sales revenue